the height of a triangle is 2cm shorter than its base. The area is 17.5 cm^2. find the base of the triangle
Answer by mananth(16949)

the height of a triangle is 2cm shorter than its base. The area is 17.5 cm^2. find the base of the triangle
let the base be x
height = x-2
Area= 1/2 * x * (x-2)
= x^2-2x /2 = = 35
x^2-2x-35=0
x^2-7x+2x-35=0
x(x-7) +2(x-7)=0
x+2)(x-7)=0
x=7 base